find the sum of the positive even integers up to and including 350.
Answer by Alan3354(69443)

find the sum of the positive even integers up to and including 350.
--------------
Even integers from 2 to 350 inclusive --> (350 - 0)/2 integers = 175 integers
----
The average is (2 + 350)/2 = 176
175*176 = 30800